OpenStudy (anonymous):

The perimeter of a rectangle is 45 feet.If the length is 12feet,find the width.

OpenStudy (anonymous):

2l+2w=perimeter so... 2*2=4 then 11-4=7, then 7divided by 2= 3.5

OpenStudy (anonymous):

anddddd thanks for the medal! lol

OpenStudy (anonymous):

your welcome